OpenStudy (anonymous):

Given the quadratic equation: 5n2 =9n2 Find all solutions for n

OpenStudy (anonymous):

\[5n^2=9n^2\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Let get them to one side

OpenStudy (anonymous):

5n^2-9n^2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

5n^2-9n^2 =0

OpenStudy (anonymous):

n^2(5-9)=0

OpenStudy (anonymous):

-4n^2=0

OpenStudy (anonymous):

so n is 0

OpenStudy (anonymous):

ok could it be n=0 or 9/5?

OpenStudy (anonymous):

no. 5(9/5)^2\[\neq\]9(9/5)^2

OpenStudy (anonymous):

do you use the same formula for this one ? Given the equation: (2x - 87)(x + 5) = 0 Find all solutions for x

OpenStudy (anonymous):

When there are in factor(parenthesis) it is actually much easier. you just look at each factor individually. First one is (2x-87) 2x=87;x=87/2 . the second factor (x+5) is just -5

OpenStudy (anonymous):

I get this one! It is easy for real!

OpenStudy (anonymous):

Thank you so much for your help!! I assume you're a teacher as you went step by step with me instead of just giving me the answer

OpenStudy (anonymous):

No, I am a student myself.I just don't think it is beneficial to give out answer.

OpenStudy (anonymous):

Thank you! I wrote the steps down and if i get the problem again It will be better for me to attempt

OpenStudy (anonymous):

what grade are you?

OpenStudy (anonymous):

Grade!!! Im in college!! Never was good with Math

OpenStudy (anonymous):

that's okay, I struggled in writing classes

OpenStudy (anonymous):

what grade are you in?

OpenStudy (anonymous):

freshman @ U of I

OpenStudy (anonymous):

U of I where is that?

OpenStudy (anonymous):

illinois

OpenStudy (anonymous):

Oh OK do you have a ffacebook account if soo look me up cally will